OpenStudy (anonymous):

A number y varies inversely as the square of x. If y = 2 when x = 6, what is y when x = 3?

OpenStudy (anonymous):

The equation is y=k/x^2 So, just plug stuff in .Do you want me to do it out for you?

OpenStudy (anonymous):

or, here: 2=k/36 Solve for K. Then plug in what you get for K, and what X is, and solve for Y.

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

Alright. So, for the first part: 2=k/36 multiple both sides by 36 in order to get rid of the fraction. this leaves you with: 72=K Got it so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Then you want to plug in K and X into the new equation. It's the SAME equation as before, so that's easy :] y=k/x^2 y=72/9 Just divide and you get: y=8 Hope that helped! Do you need any more help on this?

OpenStudy (anonymous):

no thanks! thanks so much :)

OpenStudy (anonymous):

No problem! Feel free to grab me if you need anything else, hon'.